# A General Theory of Static Response for Markov Jump Processes

### v1

@@ Abstract
Unicyclic networks.— We further restrict our theory to systems with a single cycle. Using Eqs. 2 and 9 for the unicyclic system, we have $P_{ee^{\prime}}=c_{e}B_{\gamma e^{\prime}}$ and $j_{e}=\mathcal{J}c_{e}$ where $\mathcal{J}$ is the flux and $c_{e}$ are the elements of the single cycle $\bm{c}$. Using Eq. 14, we find that if $j_{e},j_{e^{\prime}}\neq 0$ then $j_{e}=j_{e^{\prime}}$ and $\phi_{e^{\prime}}d_{p_{e^{\prime}}}\ln j_{e}=P_{e^{\prime}e^{\prime}}$, which results in the following bounds [see Eq. 15]
$0\leq\phi_{e^{\prime}}d_{p_{e^{\prime}}}\ln j_{e}\leq 1\,,$ (18)
where $e,e^{\prime}\in\mathcal{E}$ cover all combinations of the perturbing $e^{\prime}$ and observing $e$ edges. Equation 18 implies that all the scaled sensitivities are positive and must add up to one due to Eq. 11. Since their individual upper bound coincides with their sum, the saturation of one automatically suppresses the others. Unicyclic systems play an important role in understanding molecular motors and metabolic networks; see, for example, the dynein model in [39] and the biochemical switch in [26]. An example of four-state unicyclic network is shown in Fig. 1a. The histogram in Fig. 1b shows that the scaled responses (color bins) sum up to one for the random rate matrices, as predicted by Eqs. 11 and 18. One also sees that, in general, the responses of $\mathcal{J}$ to edge parameters are shared between all edges, while when one tends to saturate, the other ones are suppressed.
@@
Multicycle systems.—Since elements of $\mathbb{C}$ can always be written using $\{0,1\}$, the parametric dependence of the matrix $\mathbb{P}$ is defined by the $N_{c}N_{e}$ elements of the matrix $\mathbb{B}$ (see Eq. 9), which satisfies $\mathbb{C}\mathbb{B}\mathbb{C}=\mathbb{C}$ due to A General Theory of Static Response for Markov Jump Processes. The matrix $\mathbb{C}$ is the full column rank and $\mathbb{B}\mathbb{C}=\mathbb{I}_{c}$ is the identity matrix of size $N_{c}$. We define $\overline{\mathbb{C}}$ as the invertible submatrix of $\mathbb{C}$ and note that it is always possible to define cycles such that ¯$\overline{\mathbb{C}}=\mathbb{I}_{c}$. Therefore, we can write
$\mathbb{B}\begin{pmatrix}\mathbb{I}_{c}\\ \tilde{\mathbb{C}}\end{pmatrix}=(\overline{\mathbb{B}},\tilde{\mathbb{B}})\begin{pmatrix}\mathbb{I}_{c}\\ \tilde{\mathbb{C}}\end{pmatrix}=\mathbb{I}_{c}\,,\quad\overline{\mathbb{B}}=\mathbb{I}_{c}-\tilde{\mathbb{B}}\tilde{\mathbb{C}}\,,$ (23)
@@
and is only possible if #$\#\text{ide}=N_{e}-1$ and $N_{c}=1$ (unicyclic models) or $N=2$ (two states models).
Beyond unicyclic and two-state models, only part of the possible responses are bounded by a linear combination of the local responses. To identify which ones, using Eqs. 9 and 23, we write
@@
(~$(\tilde{\mathbb{C}}\tilde{\mathbb{B}})_{ee}$ $=P_{ee}\,,\quad e>N_{c}\,,$ (25b)
@@
Thermodynamic responses.—The rates can be expressed in terms of their symmetric $v_{e}=v_{-e}=\ln\sqrt{W_{+e}W_{-e}}$ and antisymmetric $w_{e}=\ln\sqrt{W_{+e}/W_{-e}}$ ($w_{-e}=\ln\sqrt{W_{-e}/W_{+e}}$) parts as $W_{+e}=\exp(v_{e}+w_{e})$ and $W_{-e}=\exp(v_{e}-w_{e})$. This decomposition is used in stochastic thermodynamics (i.e. for open systems undergoing transitions caused by thermal reservoirs) because the antisymmetric part defines the energetics of the system [40, 41]:
$\ln\frac{W_{+e}}{W_{-e}}=\sum_{i\in\mathcal{S}}E_{i}S_{ie}+F_{e}\,,\quad F_{e}$ $=\sum_{\alpha\in\mathcal{F}}X_{e\alpha}f_{\alpha}\,.$ (26)
The l.h.s. of Eq. 26 is the entropy change in the reservoirs caused by a transition $e$. It can always be split, in the r.h.s., as a change in the Massieu potential $E_{i}$ of the state $i$ and as a non-conservative contribution $F_{e}=-F_{-e}$. The latter can be expressed, in Eq. 26, as a sum over a subset of reservoirs $\mathcal{F}$ where each term consists of an amount of conserved quantities exchanged with the reservoirs during a transition $e$, $X_{e\alpha}$, multiplied by a conjugated (fundamental [40]) thermodynamic force, $f_{\alpha}$ made of differences of intensive fields of the reservoirs such as inverse temperatures or chemical potentials [40]. Using Eq. 26, we find the responses
$d_{f_{\alpha}}j_{e}$ $=\sum_{\rho\in\mathcal{E}}d_{f_{\alpha}}F_{\rho}d_{F_{\rho}}j_{e}=\sum_{\rho\in\mathcal{E}}X_{\rho\alpha}P_{e\rho}\partial_{F_{\rho}}j_{\rho}\;,$ (27)
which can be obtained analytically for known $\partial_{F_{\rho}}j_{\rho}$. Since the (fundamental [40]) currents exchanged with the different reservoirs are the elements of the vector $\bm{I}=\mathbb{X}^{\intercal}j$, their response to changes in the thermodynamic forces has the following form:
$\mathbb{R}^{I}=[d_{f_{\alpha^{\prime}}}I_{\alpha}]_{\{\alpha,\alpha^{\prime}\}}=\mathbb{X}^{\intercal}\mathbb{P}\mathbb{J}(\bm{F})\mathbb{X}=\mathbb{X}^{\intercal}\mathbb{R}^{j}\mathbb{X}\,,$ (28)
where $\alpha,\alpha^{\prime}\in\mathcal{F}$ and $\mathbb{J}(\bm{F})=\text{diag}(\dots,\partial_{F_{e}}j_{e},\dots)$. Equation 28 shows the linear transformation from the edge responses to the responses of the reservoir currents.
Let us assume now that $w_{e}$ is independent of $v_{e}$. This is relevant, for example, for Arrhenius-like rates [24], as well as for electron transfer rates in CMOS transistors [42] or single electron tunneling rates in the wide-band approximation [43]. Equation 26 shows that $w_{e}$ only depends on the perturbation of the energy and forces, but does not depend on the perturbation of the kinetic parameters. Such kinetic $v_{e}$ and energy (thermodynamic forces) $w_{e}$ perturbations will be constrained by Eq. 11. Calculating the partial derivatives $\partial_{w_{e}}j_{e}=\tau_{e}$ and $\partial_{v_{e}}j_{e}=j_{e}$, we find $\phi_{e}=j_{e}/\tau_{e}$ (resp. $\phi_{e}=1$) for $w_{e}$ (resp. $v_{e}$), where $\tau_{e}=W_{e}\pi_{s(+e)}+W_{-e}\pi_{s(-e)}$ is the edge traffic. Inserting $\phi_{e}$ into Eq. 11, we arrive at the Symmetric and Antisymmetric SRRs
@@
We note that unlike the antisymmetric parameterisation, the symmetric one is homogeneous $\bm{h}=(\dots,v_{e},\dots)^{\intercal}$: the symmetric RSSs in A General Theory of Static Response for Markov Jump Processes coincide with Eqs. 4 and 5.
Figure 2: (a): The graph for double QDD model [40] with $W_{\pm e}=\Gamma_{e}(1+\exp(\pm F_{e}))^{-1}$, where $\Gamma_{e}$ are the electronic tunneling rates. (b): The disks indicate the bounded scaled responses ($20$ out of $36$). (c): Illustration of the CRR for (a), where $N_{c}=3$. The heights of color bins (from purple to yellow) correspond to $\phi_{e}d_{F_{e}}\ln j_{e}$ for $e=1,\dots,6$. (d): Numerical calculations of $(\phi_{e^{\prime}}/j_{e^{\prime}})d_{F_{e^{\prime}}}j_{e}$ for bounded configurations shown in $(b)$. (e): Numerical calculations as in (c), but for unbounded configurations. Rates $W_{\pm e}$ are homogeneously distributed in the range $0\leq W_{\pm e}\leq W_{m}$, where $W_{m}=1000$.
Application.—We illustrate our approach using a Double Quantum Dots (DQD) model [40] with three cycles (see the graph in Fig. 2a). The topology of the DQD is defined by:
$\footnotesize\mathbb{S}=\begin{pmatrix}-1&-1&0&0&-1&0\\ 1&0&-1&0&0&-1\\ 0&1&0&-1&1&0\\ 0&0&1&1&0&1\end{pmatrix}\,,\overline{\mathbb{C}}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}\,,\tilde{\mathbb{C}}=\begin{pmatrix}-1&0&0\\ -1&-1&0\\ 1&0&-1\end{pmatrix}\,.$ (40)
The topology of the cycles corresponds to #$\#\text{ide}=5$. Solving the independent subsystem of Eqs. (A General Theory of Static Response for Markov Jump Processes), we arrive at
$\tilde{\mathbb{B}}=\begin{pmatrix}-P_{44}&P_{22}-P_{55}-1&P_{66}+P_{33}-1\\ \tilde{B}_{21}&P_{22}-1&\tilde{B}_{23}\\ \tilde{B}_{31}&\tilde{B}_{32}&P_{33}-1\end{pmatrix}\,.$ (44)
Calculating the matrix $\mathbb{P}=\mathbb{C}\big{(}\mathbb{I}-\tilde{\mathbb{B}}\tilde{\mathbb{C}},\tilde{\mathbb{B}}\big{)}$, we find all elements $P_{ee^{\prime}}$ which are the linear combination of diagonal elements and bounded. We illustrated them in Fig. 2(b), where more than half of the responses are bounded as $-1\leq(\phi_{e^{\prime}}/j_{e^{\prime}})d_{p_{e^{\prime}}}j_{e}\leq 1$ for $(e,e^{\prime})$ marked by disks in Fig. 2(b). We note that unlike the always postive local responses, the nonlocal ones can be negative; see numerical simulations in Fig. 2(cd).
Future studies.— Our approach provides powerful tools to identify networks that are highly sensitive or extremely resilient to perturbations. It is also ideally suited to study the responses of enzymatic changes (proofreading, sensing) in chemical reaction networks, in particular in conjunction with recently developed circuit theory [44].
Acknowledgements—This research was supported by Fonds National de la Recherche—FNR, Luxembourg: Project ChemComplex (C21/MS/16356329) and T.A. by Project ThermoElectroChem (C23/MS/18060819).
Appendix A Derivation of A General Theory of Static Response for Markov Jump Processes
The total derivative $d_{p}$ of Eq. 2 on $p\in\bm{p}$ reads
@@ Appendix A Derivation of A General Theory of Static Response for Markov Jump Processes
$\footnotesize\sum_{j\in\hat{\mathcal{S}}}\sum_{e\in\mathcal{E}}S_{ie}(\Gamma_{ej}-\Gamma_{eN})d_{p}\pi_{j}$ $=-\sum_{e\in\mathcal{E}}\sum_{k\in\mathcal{S}}S_{ie}(\partial_{p}\Gamma_{ek})\pi_{k}\,,$
$\sum_{j\in\hat{\mathcal{S}}}K_{ij}d_{p}\pi_{j}$ $=-\sum_{k\in\mathcal{S}}S_{ie}J_{ep}\,,$ (46)
where $\hat{\mathcal{S}}=\mathcal{S}\setminus\{N\}$ denotes the space of $N-1$ states, $J_{ep}$ are elements of the Jacobian defined in Eq. 7 and where we introduce the matrix
@@
We proceed with the proof that the matrix $\mathbb{K}$ is invertible. The auxiliary matrix $\mathbb{K}_{N}$ coincides with $\mathbb{W}$, except the $N$th row, which is the row of ones. We first prove that $\det\mathbb{K}=\det\mathbb{K}_{N}$ as $\det\mathbb{K}_{N}=$
@@
In the second, using the results of Ref. [24] for $\det\mathbb{K}_{N}$ we arrive at:
@@
where $\lambda_{i}$ are nonzero eigenvalues of the matrix $\mathbb{W}$. Thus, Eq. 53 implies that $\mathbb{K}$ is invertible.
Multiplying Appendix A by $\mathbb{K}^{-1}$, we find the responses $d_{p}\pi_{i}$. Using them to calculate $d_{p}\pi_{N}$ and rewriting the result in matrix form, we arrive at Eq. 6c. Similarly with Eq. 45, we calculate the response $d_{p}\bm{j}$ of the current:
@@
which has the matrix form shown in Eq. 6d.
Appendix B Independent diagonal elements
[…]
$\mathbb{C}=\text{diag}\big{(}\dots,\mathbb{C}_{i},\dots\big{)}\,,$ (55)
where $\mathbb{C}_{i}$ are full column rank matrices and all other elements except those matrices are zero. The examples in the main text have only one block $\mathbb{C}_{1}=\mathbb{C}$. The blocks in Eq. 55 correspond to the cycles without common edges; see Fig. 3. We prove that every diagonal block in Eq. 55 adds one constraint to the diagonal elements, that is, equivalent #$\#\text{ide}=N_{e}-N_{b}$, where $N_{b}$ is the number of blocks $\mathbb{C}_{i}$ (disconnected cycles in $\mathcal{G}$). Consider the block $\mathbb{C}_{i}$ of size $n_{i}\times\gamma_{i}$. We use the decomposition $\mathbb{P}=\mathbb{C}\mathbb{B}$ for block $\mathbb{C}_{i}$ as $\mathbb{P}_{i}=\mathbb{C}_{i}\mathbb{B}_{i}$, where $\mathbb{B}_{i}$ and $\mathbb{P}_{i}$ are diagonal blocks of the matrices $\mathbb{B}$ and $\mathbb{P}$, respectively. In addition, we can use the condition $\mathbb{B}\mathbb{C}=\mathbb{I}$ in block form as $\mathbb{B}_{i}\mathbb{C}_{i}=\mathbb{I}_{\gamma_{i}}$, where $\mathbb{I}_{\gamma_{i}}$ is the identity matrix of size $\gamma_{i}$. Calculating the trace of the block $\mathbb{P}_{i}$ we find
Appendix C Supplementary: Proof of Eq. 15
Here we prove that the diagonal elements $P_{ee}$ are bounded as $0\leq P_{ee}\leq 1$. We introduce the projection matrix $\mathbb{\Pi}=\mathbb{I}-\mathbb{P}$ with the diagonal elements defined by

### v2

@@ Abstract
Unicyclic networks.— We further restrict our theory to systems with a single cycle. Unicyclic systems play an important role in understanding molecular motors and metabolic networks; see e.g. the dynein model in [47] and the biochemical switch in [28]. Choosing an edge orientation such that $\bm{c}$ consists only of $1$s and 00s and using Eqs. 2 and 9 for the unicyclic system, we have $P_{ee^{\prime}}=c_{e}B_{1e^{\prime}}$ and $j_{e}=\mathcal{J}c_{e}$ where $\mathcal{J}$ is the flux and $c_{e}$ are the elements of the single cycle $\bm{c}$, that result in $P_{ee^{\prime}}=c_{e}/c_{e^{\prime}}P_{e^{\prime}e^{\prime}}=j_{e}/j_{e^{\prime}}P_{e^{\prime}e^{\prime}}$ Using Eq. 14 and that if $j_{e},j_{e^{\prime}}\neq 0$ then $j_{e}=j_{e^{\prime}}$, we find $P_{e^{\prime}e^{\prime}}=P_{ee^{\prime}}=(\partial_{p_{e^{\prime}}}j_{e^{\prime}})^{-1}d_{p_{e}}j_{e}$ which results in the following bounds [see Eq. 15]
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0\leq\Big{(}\frac{\partial j_{e^{\prime}}}{\partial p_{e^{\prime}}}\Big{)}^{-1}\frac{dj_{e}}{dp_{e^{\prime}}}\leq 1}$ (18)
far any combinations of the perturbing $e^{\prime}$ and observing $e$ edges. In addition, the SRR (11) for unicyclic network becomes
$\sum_{e=1}^{N_{e}}\Big{(}\frac{\partial j_{e}}{\partial p_{e}}\Big{)}^{-1}\frac{dj_{e^{\prime}}}{dp_{e}}=1\,.$ (19)
Since scaled responses are nonnegative and add up to one, the saturation of one automatically suppresses all the others.
Physical example: By removing the third reservoir (edge 5 and 6) and changing the orientation of the edges 2 and 4 in Fig. 1, the model becomes a four-state unicyclic network, see Fig. 2a. The histogram in Fig. 2b illustrates that the scaled responses to edge parameters $\bm{\Psi}$ are nonnegative and sum up to one as predicted by Eqs. 18 and 19. One also sees that they are typically shared between all edges, whereas when one tends to saturate, the other ones are suppressed. In Sec. D of [41] we derive the following tight bounds for the responses of any edge current $j_{e}$ to the energy levels $\epsilon_{u}$ and $\epsilon_{d}$: $-|\partial_{\Psi_{1}}j_{1}|\leq T_{1}d_{\epsilon_{u}}j_{e}\leq|\partial_{\Psi_{4}}j_{4}|$ and $-|\partial_{\Psi_{3}}j_{3}|\leq T_{2}d_{\epsilon_{d}}j_{e}\leq|\partial_{\Psi_{2}}j_{2}|$, which are illustrated for different values of the thermodynamic force $U_{C}(\beta_{1}-\beta_{2})$ in Fig. 2(c,d).
@@
Multicycle systems.—Since the elements of $\mathbb{C}$ can always be written using $\{0,1\}$, the parametric dependence of the matrix $\mathbb{P}$ is defined by the $N_{c}N_{e}$ elements of the matrix $\mathbb{B}$ in Eq. 9, which satisfies $\mathbb{C}\mathbb{B}\mathbb{C}=\mathbb{C}$ due to Eq. 8. Matrix $\mathbb{C}$ is full column rank and $\mathbb{B}\mathbb{C}=\mathbb{I}_{c}$ is the identity matrix of size $N_{c}$. Defining $\overline{\mathbb{C}}$ as the invertible submatrix of $\mathbb{C}$ and noting that it is always possible to define cycles such that ¯$\overline{\mathbb{C}}=\mathbb{I}_{c}$,
$\mathbb{B}\mathbb{C}=\mathbb{B}\begin{pmatrix}\mathbb{I}_{c}\\ \tilde{\mathbb{C}}\end{pmatrix}=(\overline{\mathbb{B}},\tilde{\mathbb{B}})\begin{pmatrix}\mathbb{I}_{c}\\ \tilde{\mathbb{C}}\end{pmatrix}=\mathbb{I}_{c}\,,\quad\overline{\mathbb{B}}=\mathbb{I}_{c}-\tilde{\mathbb{B}}\tilde{\mathbb{C}}\,,$ (20)
@@
which is only possible if #$\#\text{ide}=N_{e}-1$, and thus if $N_{c}=1$ (unicyclic models) or if $N=2$ (two states models).
Beyond unicyclic and two-state models, only part of the possible responses are bounded by a linear combination of the local responses. To identify which ones, using Eqs. 9 and 20, we write
@@
(~$(\tilde{\mathbb{C}}\tilde{\mathbb{B}})_{ee}$ $=P_{ee}\,,\quad e>N_{c}\,.$ (22b)
@@
Thermodynamic responses.—The rates can be expressed in terms of their symmetric $v_{e}\equiv\ln\sqrt{W_{+e}W_{-e}}=v_{-e}$ and antisymmetric $w_{e}\equiv\ln\sqrt{W_{+e}/W_{-e}}=-w_{-e}$ parts as $W_{\pm e}=\exp(v_{e}\pm w_{e})$. For the QDs of Fig. 1, $v_{e}=\ln\Gamma_{e}$ and $w_{e}=-\ln(1+\exp(\Psi_{e}))$. This decomposition is used in stochastic thermodynamics (i.e. for open systems undergoing transitions caused by thermal reservoirs) because the antisymmetric part defines the energetics of the system [44, 48]:
$\ln\frac{W_{+e}}{W_{-e}}=\sum_{i\in\mathcal{S}}E_{i}S_{ie}+\mathcal{F}_{e}\,,\quad\mathcal{F}_{e}$ $=\sum_{\alpha\in\mathcal{R}}X_{e\alpha}f_{\alpha}\,.$ (23)
The l.h.s. of Eq. 23 is the entropy change in the reservoirs caused by a transition $e$. It can always be split, in the r.h.s., as a change in the Massieu potential $E_{i}$ of the state $i$ and as a non-conservative contribution $\mathcal{F}_{e}=-\mathcal{F}_{-e}$. The latter can be expressed, in Eq. 23, as a sum over a subset of reservoirs $\mathcal{R}$ where each term consists of an amount of conserved quantities exchanged with the reservoirs during a transition $e$, $X_{e\alpha}$, multiplied by a conjugated (fundamental [44]) thermodynamic force, $f_{\alpha}$ made of differences of intensive fields of the reservoirs such as inverse temperatures or chemical potentials [44]. Using Eq. 23 and Eq. 14, we find
$d_{f_{\alpha}}j_{e}$ $=\sum_{\rho\in\mathcal{E}}d_{f_{\alpha}}\mathcal{F}_{\rho}d_{\mathcal{F}_{\rho}}j_{e}=\sum_{\rho\in\mathcal{E}}X_{\rho\alpha}P_{e\rho}\partial_{\mathcal{F}_{\rho}}j_{\rho}\;,$ (24)
which can be obtained analytically for known $\partial_{\Psi_{\rho}}j_{\rho}$. Since the (fundamental [44]) currents exchanged with the different reservoirs are the elements of the vector $\bm{I}=\mathbb{X}^{\intercal}j$, their response to the thermodynamic forces read
[…]
Let us now assume that $w_{e}$ is independent of $v_{e}$. This is relevant, for example, for Arrhenius-like rates [25], as well as for electron transfer rates in CMOS transistors [49] or single electron tunneling rates in the wide-band approximation [46]. Equation 23 shows that $w_{e}$ depends only on the perturbation of the energy and forces, but does not depend on the perturbation of the kinetic parameters. Such kinetic $v_{e}$ and energy (thermodynamic forces) $w_{e}$ perturbations will be constrained by Eq. 11. Calculating the partial derivatives $\partial_{w_{e}}j_{e}=\tau_{e}$ and $\partial_{v_{e}}j_{e}=j_{e}$, we find $\phi_{e}=j_{e}/\tau_{e}$ (resp. $\phi_{e}=1$) for $w_{e}$ (resp. $v_{e}$), where $\tau_{e}\equiv W_{e}\pi_{s(+e)}+W_{-e}\pi_{s(-e)}$ is the edge traffic. Inserting $\phi_{e}$ into Eq. 11, we arrive at the Symmetric and Antisymmetric SRRs
@@
We note that unlike the antisymmetric parameterisation, the symmetric one is homogeneous $\bm{h}=(\dots,v_{e},\dots)^{\intercal}$ as the symmetric RSSs in Eq. 26 coincide with Eqs. 4 and 5.
Figure 3: For the model in Fig. 1: (a) The disks indicate the $20$ bounded $P_{ee^{\prime}}$’s out of $36$. (b) Validity of the CRR [Eq. 17] with $N_{c}=3$. The heights of the color bins (from black to yellow) correspond to $P_{ee}$ for $e=1,\dots,6$ and to randomly and homogeneously distributed $0\leq\epsilon_{u},\epsilon_{d},U_{C}\leq 5$, $0<\Gamma_{e}\leq 1000$ and $-10\leq\mu_{2},\mu_{3}\leq 10$, with $T_{i}=1$, $\mu_{1}=0$. (c) $P_{ee^{\prime}}$ corresponding to the red disks in $(a)$, as a function of the thermodynamic force $\beta(\mu_{3}-\mu_{2})$. (d) Physical responses $|d_{\epsilon_{d}}j_{1}|$ in units of $R=\sum_{e\neq 1,4}|\partial_{\Psi_{e}}j_{e}|$. Dashed lines denote our bounds.
Physical example: For the QDs in Fig. 1, we use Eq. 22 to find all the elements $P_{ee^{\prime}}$ which are linear combinations of diagonal elements and are thus bounded, see Sec. G of [41] and Fig. 3a. The $6$ local scaled responses, $P_{ee}$, sum to $N_{c}=3$ as predicted by the CRR [Eq. 17], see Fig. 3b. The nonlocal scaled responses can be negative, but those marked as disks in Fig. 3a are bounded as $-1\leq P_{ee^{\prime}}\leq 1$, see Eq. (G4) in [41] and Fig. 3c. We use the properties of the matrix $\mathbb{P}$ to bound the responses of the current to physical parameters in Sec. H of [41]. We find that $d_{p}j_{\alpha}=\sum_{e}P_{\alpha e}\partial_{p}\Psi_{e}$ for $\alpha={1,4}$, where $-1\leq P_{\alpha e}\leq 1$. We also find $|d_{\epsilon_{d}}j_{\alpha}|\leq R$, where $R=|\partial_{\Psi_{2}}j_{2}|+|\partial_{\Psi_{3}}j_{3}|+|\partial_{\Psi_{5}}j_{5}|+|\partial_{\Psi_{6}}j_{6}|$. This is illustrated numerically in Fig. 3d for different values of the thermodynamic force $\beta(\mu_{3}-\mu_{2})$, where we see that large responses arise far from equilibrium.
Future studies.— Our approach provides powerful tools to identify networks that are highly sensitive or extremely resilient to perturbations. It is also ideally suited to study the responses of enzymatic changes (proofreading, sensing) in chemical reaction networks, in particular in conjunction with recently developed circuit theory [50]. Extending our approach to non-stationary response theory of Markov processes as in Refs. [51, 52, 35, 36] is also an interesting perspective.
Acknowledgements—This research was supported by Fonds National de la Recherche—FNR, Luxembourg: M. E. and T.A. by Project ChemComplex (C21/MS/16356329) and T. A. also by Project ThermoElectroChem (C23/MS/18060819).
Appendix A Derivation of Eq. 6.
Here we derive the matrix expressions (6) for the probability and current responses to an arbitrary parameterisation $\bm{p}$. The total derivative $d_{p}$ of Eq. (2) on $p\in\bm{p}$ reads
@@
$\footnotesize\sum_{j\in\hat{\mathcal{S}}}\sum_{e\in\mathcal{E}}S_{ie}(\Gamma_{ej}-\Gamma_{eN})d_{p}\pi_{j}$ $=-\sum_{e\in\mathcal{E}}\sum_{k\in\mathcal{S}}S_{ie}(\partial_{p}\Gamma_{ek})\pi_{k}\,,$ (28)
where $\hat{\mathcal{S}}=\mathcal{S}\setminus\{N\}$ denotes the space of $N-1$ states, or
$\footnotesize\sum_{j\in\hat{\mathcal{S}}}K_{ij}d_{p}\pi_{j}=-\sum_{e\in\mathcal{E}}S_{ie}J_{ep}\,,$ (29)
where $J_{ep}=\partial_{p}j_{e}$ are elements of the Jacobian defined in Eq. (7) and where we introduced the matrix
@@
We proceed with the proof that the matrix $\mathbb{K}$ is invertible. The auxiliary matrix $\mathbb{K}_{N}$ coincides with $\mathbb{W}$, except the $N$th row, which is the row of ones. Therefore $\det\mathbb{K}_{N}=$
@@
Then, using the results of Ref. [25] for $\det\mathbb{K}_{N}$, we arrive at:
@@
where $\lambda_{i}$ are nonzero eigenvalues of the matrix $\mathbb{W}$, which implies that $\mathbb{K}$ is invertible.
Multiplying Eq. 29 by $\mathbb{K}^{-1}$, we find the responses $d_{p}\pi_{i}$. Using them to calculate $d_{p}\pi_{N}$ and rewriting the result in matrix form, we arrive at Eq. (6a). Similarly, using Eq. 27 we can calculate the response of the current:
@@
which has the matrix form shown in Eq. (6b).
Appendix B Physical Parameters in Fig. 1
To calculate the Jacobian (7) for physical parameter $p$ from the model of Fig. 1, we apply the chain rule:
$\mathbb{J}$ $=\Big{[}\partial_{p}j_{e}\Big{]}_{\{e,p\}}=\Big{[}\sum_{e^{\prime}}\partial_{\Psi_{e^{\prime}}}j_{e}\partial_{p}\Psi_{e^{\prime}}\Big{]}_{\{e,p\}}=\Big{[}\partial_{\Psi_{e}}j_{e}\partial_{p}\Psi_{e}\Big{]}_{\{e,p\}}\,,$ (34)
where we used that $j_{e}$ explicitly depends on $\Psi_{e}$ only.
(i) The set $\bm{p}_{1}=\{\epsilon_{u},\epsilon_{d},U_{c}\}$ results in the Jacobian
$\mathbb{J}_{1}=\begin{pmatrix}\beta_{1}\partial_{\Psi_{1}}j_{1}&0&0\\ 0&\beta_{2}\partial_{\Psi_{2}}j_{2}&0\\ 0&\beta_{2}\partial_{\Psi_{3}}j_{3}&\beta_{2}\partial_{\Psi_{3}}j_{3}\\ \beta_{1}\partial_{\Psi_{4}}j_{4}&0&\beta_{1}\partial_{\Psi_{4}}j_{4}\\ 0&\beta_{3}\partial_{\Psi_{5}}j_{5}&0\\ 0&\beta_{3}\partial_{\Psi_{6}}j_{6}&\beta_{3}\partial_{\Psi_{6}}j_{6}\end{pmatrix}\,,$ (35)
which has the rank $\text{rk}~{}\mathbb{J}_{1}=3$. Since $\text{rk}~{}\mathbb{J}_{1}<N_{e}=6$, the set $\bm{p}_{1}$ does not satisfy the conditions of the response relations [Eq. 10].
(ii) The Jacobing of the set $\bm{p}_{2}=\{\epsilon_{u},\epsilon_{d},U_{c},\mu_{1},\mu_{2},\mu_{3}\}$ reads
$\mathbb{J}_{2}=\begin{pmatrix}\beta_{1}\partial_{\Psi_{1}}j_{1}&0&0&-\beta_{1}\partial_{\Psi_{1}}j_{1}&0&0\\ 0&\beta_{2}\partial_{\Psi_{2}}j_{2}&0&0&-\beta_{2}\partial_{\Psi_{2}}j_{2}&0\\ 0&\beta_{2}\partial_{\Psi_{3}}j_{3}&\beta_{2}\partial_{\Psi_{3}}j_{3}&0&-\beta_{2}\partial_{\Psi_{3}}j_{3}&0\\ \beta_{1}\partial_{\Psi_{4}}j_{4}&0&\beta_{1}\partial_{\Psi_{4}}j_{4}&-\beta_{1}\partial_{\Psi_{4}}j_{4}&0&0\\ 0&\beta_{3}\partial_{\Psi_{5}}j_{5}&0&0&0&-\beta_{3}\partial_{\Psi_{5}}j_{5}\\ 0&\beta_{3}\partial_{\Psi_{6}}j_{6}&\beta_{3}\partial_{\Psi_{6}}j_{6}&0&0&-\beta_{3}\partial_{\Psi_{6}}j_{6}\end{pmatrix}\,,$ (36)
where $N_{p}=N_{e}$, but $\text{rk}~{}\mathbb{J}_{2}=4<N_{e}=6$ means that $\bm{p}_{2}$ does not work for the response relations.
(iii) For the set $\bm{p}_{2}=\{\epsilon_{u},\epsilon_{d},U_{c},\beta_{1},\beta_{2},\beta_{3}\}$ we have
$\mathbb{J}_{3}=\begin{pmatrix}\beta_{1}\partial_{\Psi_{1}}j_{1}&0&0&(\epsilon_{u}-\mu_{1})\partial_{\Psi_{1}}j_{1}&0&0\\ 0&\beta_{2}\partial_{\Psi_{2}}j_{2}&0&0&(\epsilon_{d}-\mu_{2})\partial_{\Psi_{2}}j_{2}&0\\ 0&\beta_{2}\partial_{\Psi_{3}}j_{3}&\beta_{2}\partial_{\Psi_{3}}j_{3}&0&(\epsilon_{d}+U_{C}-\mu_{2})\partial_{\Psi_{3}}j_{3}&0\\ \beta_{1}\partial_{\Psi_{4}}j_{4}&0&\beta_{1}\partial_{\Psi_{4}}j_{4}&(\epsilon_{u}+U_{C}-\mu_{1})\partial_{\Psi_{4}}j_{4}&0&0\\ 0&\beta_{3}\partial_{\Psi_{5}}j_{5}&0&0&0&-(\epsilon_{d}-\mu_{3})\partial_{\Psi_{5}}j_{5}\\ 0&\beta_{3}\partial_{\Psi_{6}}j_{6}&\beta_{3}\partial_{\Psi_{6}}j_{6}&0&0&(\epsilon_{d}+U_{C}-\mu_{3})\partial_{\Psi_{6}}j_{6}\end{pmatrix}\,,$ (37)
which satisfies the conditions of summation relations with $\text{rk}~{}\mathbb{J}_{3}=6=N_{e}$. In addition for $\bm{p}_{2}$ the following matrix is nonsingular:
$\det\big{(}[\partial_{p_{e^{\prime}}}\Psi_{e}]_{\{e,e^{\prime}\}}\big{)}\neq 0\,,$ (38)
which implies that controlling $d\bm{p}$ one can change $d\bm{\Psi}$ independently at every edge. Thus, for Fig. 1 it is possible to use $\bm{\Psi}$ as an edge perturbation. Since $\mathbb{J}(\bm{\Psi})$ and $\mathbb{J}(\bm{\Gamma})$ are diagonal, the two scaled responses are identical $P_{ee^{\prime}}=(\partial_{\Gamma_{e^{\prime}}}j_{e^{\prime}})^{-1}d_{\Gamma_{e^{\prime}}}j_{e}=(\partial_{\Psi_{e^{\prime}}}j_{e^{\prime}})^{-1}d_{\Psi_{e^{\prime}}}j_{e}$.
Appendix C Proof of Eq. (15)
Here we prove the bounds in Eq. (15) of the main text. We introduce the projection matrix $\mathbb{\Pi}=\mathbb{I}-\mathbb{P}$ with the diagonal elements defined by